# The Male CEO and the Female Assistant: Evaluation and Mitigation of Gender Biases in Text-To-Image Generation of Dual Subjects

### v3

@@ Abstract
Recent large-scale T2I models like DALLE-3 have made progress in reducing gender stereotypes when generating single-person images. However, significant biases remain when generating images with more than one person. To systematically evaluate this, we propose the Paired Stereotype Test (PST) framework, which queries T2I models to depict two individuals assigned with male-stereotyped and female-stereotyped social identities, respectively (e.g. “a CEO” and “an Assistant”). This contrastive setting often triggers T2I models to generate gender-stereotyped images. Using PST, we evaluate two aspects of gender biases – the well-known bias in gendered occupation and a novel aspect: bias in organizational power. Experiments show that over 74% images generated by DALLE-3 display gender-occupational biases. Additionally, compared to single-person settings, DALLE-3 is more likely to perpetuate male-associated stereotypes under PST. We further propose FairCritic, a novel and interpretable framework that leverages an LLM-based critic model to i) detect bias in generated images, and ii) adaptively provide feedback to T2I models for improving fairness. FairCritic achieves near-perfect fairness on PST, overcoming the limitations of previous prompt-based intervention approaches.
@@
Yixin Wan  and Kai-Wei Chang Department of Computer Science University of California, Los Angeles {elaine1wan, kwchang}@cs.ucla.edu
@@ 3.1 Bias Definition
We formally define gender biases in T2I models as the over-generalized association between social attributes and specific gender traits in image generation. Such social attributes can manifest in a variety of forms, such as professions, social status, etc, resulting in multifaceted and intricate biases. We mainly seek to investigate Biases in Gendered Occupation and Biases in Organizational Power. Ideally, a fair model should be equally likely to generate individuals with either gender traits, regardless of their occupation or power levels.
@@ 3.2.1 Task Formulation
PST prompts T2I models to simultaneously generate one individual with male-stereotypical social attributes and another with female-stereotypical attributes. For instance, given that the occupation “CEO” is stereotypically associated with males and “assistant” with females, PST formulates a setting to prompt models to generate a CEO and an assistant in the same image. Intuitively, this contrastive dual-subject-based setting is challenging for T2I models—models easily adhere to both stereotypes for the opposite genders, e.g. generating a male CEO and a female assistant.
@@ 3.2.3 Evaluation Metrics
$Stereo(g_{img},g_{stereo})=\left\{\begin{array}[]{l}1,\;\text{if }g_{img}=g_{stereo}\\ -1,\;\text{otherwise}\\ \end{array}\right.\vspace{-0.2em}$
@@ 4.1 Model Selection
We explore gender biases in OpenAI’s DALLE-3 model and an off-the-shelf Stable Diffusion (SD) model  [^4] https://huggingface.co/stabilityai/stable-diffusion-2 with default hyperparameter settings. Since the SD model fails to generate most images with two individuals in the paired scenario, we finally choose DALLE-3 for further bias evaluation experiments. Appendix C demonstrates a few failure cases in SD’s generations.
@@ 4.4.2 Biases in Gendered Occupation
Occupation-Level SS Figure 8 visualizes the occupation-level stratified SS for biases in gendered occupation. While noticeable levels of gender biases are found in all occupations, images generated for professions such as “sheriff”, “mechanician”, and “secretary” tend to carry the highest level of biases. Full quantitative results of the stratified SS for gendered occupation biases are in Appendix F.2, Table 11.
@@ 4.5 Biases in Organizational Power
We explore how gender biases in stereotypical organizational power dynamics are unveiled under the PST setting. Figure 9 shows qualitative examples of bias manifestation in PST, compared with single-subject generation. Appendix E, Figure 12 provides additional illustrations.
@@
Figure 9: Examples of gender biases manifestation in organizational power under PST, despite the seemingly unbiased single-persona image generations. Remarkable fairness improvement can also be observed after applying our FairCritic method. The green “-” sign indicates anti-stereotype, whereas the red “+” sign highlights gender-stereotyped images.
We found that compared to the single-person generation setting, model generations in PST demonstrate a considerably greater adherence to male stereotypes across both bias aspects. [Elaine Wan: short]
@@ 5.1 FairCritic for Bias Mitigation
Experiment Setup We conducted experiments to prove the effectiveness of our proposed FairCritic approach. Since the bias judgement step requires multi-image understanding, we use gpt-4-vision-preview as the critic model. While the critic feedback framework can autonomously run for multiple loops, we limit our experiments to contain at most 1 loop due to computational constraints.
Experiment Results Results in Table 1 prove the effectiveness of the proposed FairCritic framework in mitigating biases under PST. Across both bias aspects, FairCritic is able to remarkably reduce bias levels to close to 0, achieving outstanding fairness performance. The method controllably reduces biases for both male- and female-stereotypical occupations, while at the same time maintaining interpretability through the intermediate critic model feedback step. For instance, in Figure 10, we see an example of the critic model providing constructive guidelines corresponding to observed biases, such as improving “role diversity”. Besides quantitative results, qualitative examples in Figure 3 and Figure 9 also show straightforward fairness improvements in generated images after applying FairCritic.
@@ 6 Discussion and Conclusion
We identify the major limitations of our study. Firstly, experiments in this work is limited to the English language. Secondly, since there lack previous literature on gender biases towards other gender minority groups (for instance, the LGBTQ community) in T2I models, we only considered the binary gender for bias evaluation. Future works should extend the scope of our analysis to include more gender groups to further probe social biases in T2I systems. Thirdly, due to the limited accessibility of Google’s Imagen model and the poor generation ability of the Stable Diffusion model, we were only able to conduct experiments on the DALLE-3 models. Extending analyses to additional T2I models is an important next step that future researchers in this direction shall explore. Lastly, due to cost constraints for human annotation, we were not able to further scale up our analysis. We call for the AI fairness research community to collectively investigate biases in T2I systems on a broader scale, to identify other aspects of biases that might persist in these models.
@@
Experiments in this study involve using large T2I generative models that were pre-trained on a myriad of internet-scraped images that already contain stereotypes. Previous studies have shown that these models might propagate or even amplify biases in their training data, but their evaluation methods are limited and over-simplified. In our work, we propose a novel evaluation setting to probe for implicit gender biases in T2I models under more challenging image generation settings. Results in our study uncover severe underlying biases of DALLE-3 in two aspects: gendered occupation and organizational power. Through our research, we hope to draw attention to the impeding fairness issues in multimodal generative systems, and warn against the direct usage of such models without scrutiny. Future researchers shall further contribute to defining, evaluating, and resolving such biases to promote safe and ethical usage of these generative tools.
@@ A.2 Biases in Organizational Power
Table 4: Lists of corporational roles with gender-stereotypical level of organizational power.
@@ Appendix D Automated Gender Classifiers
[Elaine Wan: add in fairface] Since the bias analysis in this work heavily relies on human-annotated gender of the individuals in generated images, we hope to explore potential options to automate the image labeling process. Previous works Bianchi et al. (2023); Seshadri et al. (2023) explored the use of Contrastive Language-Image Pre-training (CLIP) models Radford et al. (2021) to classify genders of individuals in images generated under the single-person setting. Since the proposed PST setting is based on a more challenging generation task, where two individuals are generated in one single image, CLIP is naturally unsuitable for gender annotation. Instead, we explore the use of 2 alternative automated approaches: (1) an off-the-shelf Bootstrapping Language-Image Pre-training (BLIP) model Li et al. (2022)666released under BSD 2-Clause License, a Visual Question Answering model, as an automated zero-shot gender classifier. (2) the widely-adopted FairFace (Kärkkäinen and Joo, 2019) demographic classifier.
@@
Appendix E Bias in Organizational Power: Additional Qualitative Examples
Figure 12 shows additional qualitative examples of gender bias manifestation in PST setting. Furthermore, we can observe the outstanding fairness performance of the FaiarCritic approach in reducing biases, even in images generated under the challenging dual-subject setting.
Appendix F Full Experimental Results
F.1 Overall Biases
@@
F.2 Biases in Gendered Occupation
@@
F.3 Biases in Organizational Power

### v4

@@ Abstract
Recent large-scale T2I models like DALLE-3 have made progress in reducing gender stereotypes when generating single-person images. However, significant biases remain when generating images with more than one person. To systematically evaluate this, we propose the Paired Stereotype Test (PST) framework, which queries T2I models to depict two individuals assigned with male-stereotyped and female-stereotyped social identities, respectively (e.g. “a CEO” and “an Assistant”). This contrastive setting often triggers T2I models to generate gender-stereotyped images. Using PST, we evaluate two aspects of gender biases – the well-known bias in gendered occupation and a novel aspect: bias in organizational power. Experiments show that over 74% images generated by DALLE-3 display gender-occupational biases. Additionally, compared to single-person settings, DALLE-3 is more likely to perpetuate male-associated stereotypes under PST. We further propose FairCritic, a novel and interpretable framework that leverages an LLM-based critic model to i) detect bias in generated images, and ii) adaptively provide feedback to T2I models for improving fairness. FairCritic achieves near-perfect fairness on PST, overcoming the limitations of previous prompt-based intervention approaches. We release our code and evaluation data at: https://github.com/elainew728/pst-fairness.
@@
Yixin Wan  and Kai-Wei Chang University of California, Los Angeles {elaine1wan, kwchang}@cs.ucla.edu
@@ 3.1 Bias Definition
We formally define gender biases in T2I models as the over-generalized association between social attributes and specific gender traits in image generation. Such social attributes can manifest in a variety of forms, such as professions, social status, etc., resulting in multifaceted and intricate biases. We mainly seek to investigate two bias aspects: Biases in Gendered Occupation and Biases in Organizational Power. Ideally, a fair model should be equally likely to generate individuals with either gender traits, regardless of their occupation or power levels.
@@ 3.2.1 Task Formulation
PST prompts T2I models to simultaneously generate one individual with male-stereotypical social attributes and another with female-stereotypical attributes. For instance, given that the occupation “CEO” is stereotypically associated with males and “assistant” with females, PST formulates a setting to prompt models to generate a CEO and an assistant in the same image. Intuitively, this contrastive dual-subject-based setting is challenging for T2I models—models easily adhere to both stereotypes for the opposite genders, e.g. generating a male CEO and a female assistant. Note that we were not able to extend our experiments to more complex settings like depicting 3+ subjects due to the limited capability of current T2I models to follow such complicated instructions, as illustrated in Appendix C.2. Nevertheless, we showed in Figure 10 that severe gender bias persists in generated cases.
@@ 3.2.3 Evaluation Metrics
$Stereo(g_{img},g_{stereo})=\left\{\begin{array}[]{l}1,\;\text{if }g_{img}=g_{stereo}\\ -1,\;\text{otherwise}\\ \end{array}\right.$
@@ 4.1 Model Selection
We explore gender biases in OpenAI’s DALLE-3 model and an off-the-shelf Stable Diffusion (SD) model  [^4] https://huggingface.co/stabilityai/stable-diffusion-2 with default hyperparameter settings. Since the SD model fails to generate most images with two individuals in the paired scenario, we finally choose DALLE-3 for further bias evaluation experiments. Appendix C.1 demonstrates a few failure cases in SD’s generations.
@@ 4.4.2 Biases in Gendered Occupation
Occupation-Level SS Figure 8 visualizes the occupation-level stratified SS for biases in gendered occupation. While noticeable levels of gender biases are found in all occupations, images generated for professions such as “sheriff”, “mechanician”, and “secretary” tend to carry the highest level of biases. Full quantitative results of the stratified SS for gendered occupation biases are in Appendix E.2, Table 11.
@@ 4.5 Biases in Organizational Power
We explore how gender biases in stereotypical organizational power dynamics are unveiled under the PST setting. Figure LABEL:examples_3 shows qualitative examples of bias manifestation in PST, compared with single-subject generation.
@@
We found that compared to the single-person generation setting, model generations in PST demonstrate a considerably greater adherence to male stereotypes across both bias aspects.
@@ 5.1 FairCritic for Bias Mitigation
Experiment Setup
@@
We conducted experiments to prove the effectiveness of our proposed FairCritic approach. Since the bias judgement step requires multi-image understanding, we use gpt-4-vision-preview as the critic model. While the critic feedback framework can autonomously run for multiple loops, we limit our experiments to contain at most 1 loop due to computational constraints.
Experiment Results Results in Table 1 prove the effectiveness of the proposed FairCritic framework in mitigating biases under PST. Across both bias aspects, FairCritic is able to remarkably reduce bias levels to close to 0, achieving outstanding fairness performance. The method controllably reduces biases for both male- and female-stereotypical occupations, while at the same time maintaining interpretability through the intermediate critic model feedback step. For instance, in Figure 9, we see an example of the critic model providing constructive guidelines corresponding to observed biases, such as improving “role diversity”. Besides quantitative results, qualitative examples in Figure 3 and Figure LABEL:examples_3 also show straightforward fairness improvements in generated images after applying FairCritic.
@@ 6 Discussion and Conclusion
Acknowledgements
We thank the UCLA-NLP+ members, conference reviewers, and conference chairs for their invaluable feedback. We also hope to acknowledge NSF #2331966, ONR grant N00014-23-1-2780, and Apple Research Award for supporting this work.
@@ Acknowledgements
We identify the major limitations of our study. Firstly, experiments in this work are limited to the English language. Secondly, since there lack previous literature on gender biases towards other gender minority groups (for instance, the LGBTQ community) in T2I models, we only considered the binary gender for bias evaluation. Future works should extend the scope of our analysis to include more gender groups to further probe social biases in T2I systems. Thirdly, due to the limited accessibility of Google’s Imagen model and the poor generation ability of the Stable Diffusion model, we were only able to conduct experiments on the DALLE-3 models. Extending analyses to additional T2I models is an important next step that future researchers in this direction shall explore. Lastly, due to cost constraints for human annotation, we were not able to further scale up our analysis. We call for the AI fairness research community to collectively investigate biases in T2I systems on a broader scale, to identify other aspects of biases that might persist in these models.
@@
Experiments in this study involve using large T2I generative models that were pre-trained on a myriad of internet-scraped images that already contain stereotypes. Previous studies have shown that these models might propagate or even amplify biases in their training data, but their evaluation methods are limited and over-simplified. In our work, we propose a novel evaluation setting to probe for implicit gender biases in T2I models under more challenging image generation settings. Results in our study uncover severe underlying biases of DALLE-3 in two aspects: gendered occupation and organizational power. Through our research, we hope to draw attention to the impending fairness issues in multimodal generative systems, and warn against the direct usage of such models without scrutiny. Future researchers shall further contribute to defining, evaluating, and resolving such biases to promote safe and ethical usage of these generative tools.
@@ A.2 Biases in Organizational Power
Table 4: Lists of corporate roles with gender-stereotypical levels of organizational power.
@@ B.1 Human Annotator Information
Figure 10: Examples of failure cases and biased cases generated by DALL-E 3 when approaching a more complex task of generating 3 people. The model fails to follow the demographic and positional information provided in the instruction; in rare cases that it successfully depicts the correct subjects, we show that biases persist.
@@ C.1 Stable Diffusion 3
C.2 DALL-E 3
Figure 10 shows failure cases in DALL-E 3 generations, in which the model either demonstrates biases or fails to generate an accurate image that is faithful to the demographic and positional information provided in the prompts. The images are generated such that in the prompts, the 3 occupations that should depicted on the left, middle, and right are indicated in order, for instance: "Generate a picture of one {} on the left, one {} in the middle, and one {} on the right." Failure cases indicate an interesting problem that current T2I models still lack the ability to follow prompts that specifically control the number of subjects generated. This aligns with recent works’ observations on T2I models’ failure on compositional T2I generation (Huang et al., 2023; Wan and Chang, 2025).
In a small proportion of cases where DALLE-3 successfully depicts 3 people, we noticed that biases persist in the 3-people generation outcomes. We demonstrate these bias cases in Figure 10 to show that T2I models continue to possess biases in multi-people generation settings beyond the paired generation setting, which again highlights the importance of our work for systematically evaluating and mitigating the issue.
@@ Appendix D Automated Gender Classifiers
Since the bias analysis in this work heavily relies on human-annotated gender of the individuals in generated images, we hope to explore potential options to automate the image labeling process. Previous works Bianchi et al. (2023); Seshadri et al. (2023) explored the use of Contrastive Language-Image Pre-training (CLIP) models Radford et al. (2021) to classify genders of individuals in images generated under the single-person setting. Since the proposed PST setting is based on a more challenging generation task, where two individuals are generated in one single image, CLIP is naturally unsuitable for gender annotation. Instead, we explore the use of 2 alternative automated approaches: (1) an off-the-shelf Bootstrapping Language-Image Pre-training (BLIP) model Li et al. (2022)666released under BSD 2-Clause License, a Visual Question Answering model, as an automated zero-shot gender classifier. (2) the widely-adopted FairFace (Kärkkäinen and Joo, 2019) demographic classifier.
@@
Appendix E Full Experimental Results
E.1 Overall Biases
@@
E.2 Biases in Gendered Occupation
@@
E.3 Biases in Organizational Power